National Inc. manufactures two models of CMD that can be used as cell phones, MPX, and digital camcorders.

Model Annual Sales
in Units
High F 11,100
Great P 17,100

National uses a volume-based costing system to apply factory overhead based on direct labor dollars. The unit prime costs of each product were as follows:

High F Great P
Direct materials $ 39.10 $ 26.50
Direct labor $ 18.70 $ 14.30
Budget factory overhead:
Engineering and Design 2,620 engineering hours $ 445,400
Quality Control 13,050 inspection hours 336,700
Machinery 33,930 machine hours 608,940
Miscellaneous Overhead 26,610 direct labor hours 132,550
Total $ 1,523,590

National's controller had been researching activity-based costing and decided to switch to it. A special study determined National's two products have the following budgeted activities:

High F Great P
Engineering and design hours 1,070 1,550
Quality control inspection hours 5,750 7,300
Machine hours 20,390 13,540
Labor hours 12,110 14,500

Using activity-based costing, applied engineering and design factory overhead for the High F model per unit is: (Rounded to the nearest cent.)

Multiple Choice

  • $5.43.

  • $13.37.

  • $16.39.

  • $32.97.

  • $68.15.
Answer #1

Answer : $ 16.39

Under Activity Based Costing :

Applied engineering overhead Rate

= Estimated Engineering and Design / Estimated Engineering Hours

= 445,400/ 2,620 = 170

for High F

Engineering overheads applied = Rate * Engineering Hours

= 1070*170 = 181900

Per Unit Engineering Cost = 181900/11,100= $16.39 (Answer)